the sum of ages of a man peter, his wife, sarah and his son is 100,twice the age of the son is sarah,the man is 5 years older than his wife, find their ages.
Answer by ankor@dixie-net.com(22740)

let m = man's age
let s = son's age
let w = wife's age
:
write an equation for each statement
:
the sum of ages of a man peter, his wife, sarah and his son is 100,
m + s + w = 100
:
twice the age of the son is sarah,
2s = w
or divide both sides by 2 and you have:
s = .5w
:
the man is 5 years older than his wife,
m + w + 5
:
Using the first equation, replace m with (w+5); replace s with .5w
(w+5) + .5w + w = 100
2.5w = 100 - 5
2.5w = 95
w = 95/2.5
w = 38 yrs is Sarah's age
then
.5(38) = 19 yr is Peter's age
and
38+5 = 43 yrs is the man's age
:
;
See if that checks out
43 + 19 + 38 = 100